What will the curve xy-3x-2y-10=0 represent?

Grade:12

2 Answers

Vikas TU
7 years ago
 xy-3x-2y-10=0Take y common,y(x – 2) = 3x + 10y = (3x + 10)/(x-2)Point of discontinuity is at x = 2It represents a curve of which relation. in y and x with other than the ellipse, parabola, hyperbola , straight line and circle.
SREEKANTH
7 years ago
xy-3x-2y-10=0Take y common,y(x – 2) = 3x + 10y = (3x + 10)/(x-2)Point of discontinuity is at x = 2It represents a curve of which relation. in y and x with other than the ellipse, parabola, hyperbola , straight line and circle.